X-ray pulses are emitted when a clump of ionized matter orbits a black hole with a period of 3.2 s. If the clump is in circular orbit, and the mass of the black hole is 23MSun, what is the orbital radius (in m)?

____ m
Answer #1

Here ,

let the radius is r

Time period, T = 3.2 s

as T^2/r^2 = 4pi^2/(G * M)

3.2^2/r^3 = 4pi^2/(6.673 *10^-11 * 23 * 2 *10^30)

solving for r

r = 9.26 *10^6 m

the radius of orbit is 9.26 *106 m